OpenStudy (anonymous):

Two cars start moving from the same point. one travels south at 10 mi/h and the other travels west at 70 mi/h. At what rate is the distance between the cars increasing 2 hours later?

hero (hero):

Take derivative with respect to time

OpenStudy (anonymous):

take the deriative of x^2 +y^2=z^2?

hero (hero):

yes

hero (hero):

What is x, y and z after two hours?

OpenStudy (anonymous):

80?

hero (hero):

x = 20 after two hours y = 140 after two hours z = ? after two hours

hero (hero):

you should be good from here. I'm sure you know how to take derviative.

hero (hero):

After that, it's just plug and play

hero (hero):

If not, I can post solution

OpenStudy (anonymous):

okay so i got 2x+2y=2z and then i plugged everything in and still got 80 lol

hero (hero):

you didn't take derviative with respect to t

hero (hero):

What you should have gotten was this: \[2x \frac{dx}{dt} +2y \frac{dx}{dt} = 2z \frac{dz}{dt} \]

hero (hero):

and of course dx/dt = 10 and dy/dt = 70

hero (hero):

you should already know x, y, and z

hero (hero):

By the way, you're solving for dz/dt